# A Guide to Tracking Phylogenies in Parallel and Distributed Agent-based Evolution Models

### v1

@@ A Practicioner’s Guide to Hereditary Stratigraphy
In simulations employing decentralized parallel and distributed computation, a reconstruction-based approach such as hereditary stratigraphy is more likely to be appropriate. Compared to perfect tracking in this setting, hereditary stratigraphy provides simpler implementation, lower runtime communication overhead, and greater robustness to data loss. Note that if annotation size is not a concern or a low generation count is anticipated, essentially perfect-quality reconstruction can be achieved with hereditary stratigraphy methods. In this case, one could simply use a retention policy that discards no differentia and a wide enough differentia width to effectively guarantee collisions will not occur (e.g., 32 or 64 bits). However, in most cases, a compromise between phylogeny approximation and per-genome annotation size will be necessary. We cover this in the next section.
@@
In most scenarios, a constant-size annotation will give better runtime performance and ensure fuller use of available memory resources. However, if you need hard guarantees on absolute or recency-proportional inference quality, an annotation size that scales $\mathcal{O}(n)$ or $\mathcal{O}(log(n))$, respectively, with generational depth will be necessary. The hstrat Python packge provides fixed_resolution_algo and recency_proportional_resolution_algo policies for such cases (moreno2022hstrat).
@@
Annotation size must compromise between memory-use and communication-bandwidth overhead and quality of reconstructed phylogenies. In cases where annotation size is not a limiting factor, using 256-bit annotation buffers with single-bit differentia will discern phylogenetic events with about 13% recency-relative precision (tilted) or 1% depth-relative precision (steady) through 1 billion generations. For full-byte differentia, discussed below, an annotation size on the order of kilobits would be very robust. Where memory use is a limiting factor, however, A 64-bit annotation buffer with single-bit differentia can give good results.
In picking annotation size, some consideration should be given to the phylogenetic scale of experimental use cases. Triplet distance (accuracy) appears to be largely stable under surveyed increases in population size and number of taxa sampled for reconstruction However, loss of inner node loss can increase when increasing reconstruction sample size. Where this is a concern, annotation size may need to be increased commensurate to intended phylogeny tip count.
Current implementations of surface algorithms are limited to buffer sizes that are even powers of two (i.e., 32, 64, 128, etc.). Where finer gradations are desired, one possibility would be to consider intermediate differentia sizes. For instance, storing 3-bit differentia over 32 surface sites would occupy 96 bits. Alternately, alternating depositions across a collection of surfaces might be considered (e.g., a 32-bit tilted surface and a 16-bit steady surface). Note, though, that the latter option would require implementation customizations akin to those used to create the hybrid surface retention policy (moreno2024hsurf). Column algorithms are more flexible in buffer sizing, but in the case of tilted retention, they make less full use of available buffer space.
@@
Hybrid retention policy is a good choice where phylodiversity-enhancing factors (ecology, spatial structure, relaxed selection pressure) are expected or expectation is unclear. Across use cases, hybrid retention often obtains competitive reconstruction quality to the better of steady and tilted policy. Even where it is outperformed by steady or tilted retention, hybrid retention avoids catastrophic failure modes characteristic, in particular, of steady retention. On platforms where steady policy implementation (necessary as a component of hybrid policy) is not readily available, tilted policy should typically suffice. If you do not expect strong phylodiversity-enhancing factors (i.e., no ecology, no spatial structure, high selection pressure), tilted policy can reliably outperform hybrid policy. In rare cases where phylodiversity-enhancing factors are very strong (e.g., pure drift conditions), a steady policy may be appropriate.
@@
Suggested default choice: use bit-size differentia.
Bit-size differentia maximize the fraction of correct reconstruction outcomes but can also introduce incorrect reconstruction outcomes. Byte-size differentia have very low incorrect reconstruction outcome rates and can reconstruct true polytomies but have a larger incidence of unresolved reconstruction outcomes (artifactual polytomies). This trade-off results in bit-size differentia giving more informative reconstructions, as measured by strict triplet distance error. If you need very strong guarantees against incorrect reconstruction outcomes, an even larger differentia size (32 or even 64 bits) may be appropriate.
@@
Another factor that might influence this decision is the software platform being designed for. Where packages providing hereditary stratigraphy are not available, surface algorithms are easier to implement owing to only needing to implement one update operation: site selection on the fixed-size surface buffer.
@@
In scenarios where explicitly differentiating between founding clades is paramount, consider adding a systematically assigned founder ID or randomly generated fixed tag. This tag would then be used as a first pass to divvy end-state annotations into independent sets before feeding them into separate reconstruction processes. If several founding lineages persist, when using single-bit differentia, spurious collisions can make completely independent clades falsely appear to share an amount of common history.
@@
It is possible to “annotate” differentia with information about genetic or phenotypic traits associated with corresponding ancestors. Conveniently, because every internal node in a hereditary stratigraphy reconstruction corresponds to a retained differentia, this approach ensures all internal nodes in a reconstructed phylogeny can be associated with a trait value. However, a similar result can be had by saving out sample “fossil” specimens (with corresponding trait information) over the course of a simulation. Because, like extant taxa, these fossils have hereditary stratigraphy annotations, they can readily be incorporated into phylogenetic reconstruction and provide information on ancestral trait states. This approach avoids bloating annotation sizes with dozens of trait values on each genome.
@@
Software implementing column-based approaches is available for Python (moreno2022hstrat) and surface-based approaches are available for Python (moreno2024hsurf) and Zig/Cerebras Software Language (moreno2024wse). Based on community feedback, C/C++ and Rust are priority targets for surface-based implementation ports. However, porting core surface algorithms to another software language should be doable with moderate effort. Surface algorithms are implemented via small, pure functions (<30$\approx$<30≈ < 30 LOC) with accompanying tests, making them amenable to stepping-stone transliteration. We found LLM assistance to be highly effective in translating the surface algorithms from Python to Zig. That said, we encourage interested researchers to reach out if hereditary stratigraphy implementation is not available in their chosen programming language — we would be happy to furnish these algorithms in other languages if you have a use case.
@@
For surface-based annotations, two data components are required: a fixed-buffer differentia store and a generation counter. The differentia store can be implemented as an array of integer data types (e.g., uint8), but for bit differentia you will likely want to use a raw memory array or an abstraction around it if available (e.g., std::bitset). Differentia store memory should be randomized upon initialization.
@@
When annotations are inherited, they should be updated to add a new differentia and elapse the annotation-associated generation counter. Usually, this should be accomplished by adding logic into an existing create offspring routine. For genotype-level tracking (as opposed to individual-level tracking), annotation updates could instead be conducted as a part of an apply mutation routine.
@@
For most applications, annotation updates (i.e., appending a new differentia) should correspond directly to generations elapsed. However, in cases where fine-resolution visibility below a certain level is not useful, annotations may be updated only that often. For tilted policy, in particular, this approach can help preserve ancient differentia coverage for longer.
In circumstances requiring time-indexed resolution or ultrametricity (tips sharing equal branch lengths from root) is necessary, annotations can be updated on the basis of logical simulation time or even real-time rather than generations elapsed. In this case, annotation updates would be elapsed, as necessary, on parents’ annotations to bring them up-to-date with the current simulation time whenever a reproduction event occurs. Such an approach may require multiple back-to-back updates if several simulation time clock cycles have elapsed. An alternate approach to achieve a time-calibrated tree would be to attach simulation timestamps with annotation differentia or to record time-stamped “fossil” taxa — as discussed above.
@@
A typical approach will sample end-state extant agents as taxa for phylogenetic reconstruction. Note that it is not necessary to exhaustively collect the entire end-state population; for many analysis use cases, a representative sample of population members will suffice. Such an approach, however, will not provide information on extinct lineages. Because hereditary stratigraphy reconstruction can stitch together taxa from across widely varying time points, specimens can be sampled during the course of an evolutionary run and be integrated directly into phylogenetic reconstruction with end-state extant genomes. Such intermediate taxa from earlier time points can serve a role akin to “fossils” in natural history. For more advanced analyses, it may be desirable to associate sampled specimens with genetic and/or phenotypic trait data (dolson2019modes; khabbazian2016fast).
@@
To export annotation data for analysis, you will need to save it to file. Workflows that save annotations separate from other genome components using a conventional plain text format (e.g., JSON, CSV) are typically most convenient. The hstrat library provides a number of plug-and-play utilities for serialization of HereditaryStratigraphicColumnObjects. Functionality to provide direct support for serializing surface data is a priority item on the hstrat road map. In the meantime, this must be done manually. For each annotation, you will need to store the generation counter and the differentia data. Hex strings provide one potentially convenient approach to encode differentia data. When loading annotation data, you will need to know the retention policy and differentia width used — so make sure these are recorded.
@@ Conclusion
In this work, we have applied empirical annotate-and-reconstruct experiments to benchmark reconstruction quality of hereditary stratigraphy approaches across use cases varying in phylogenetic structure, scale, and allocated annotation space. In these experiments, we consider,
  • differentia retention, whether annotation space should be allocated for finer resolution in discerning recent phylogenetic events;
  • annotation implementation, comparing existing column-based approaches to newer surface-based approaches optimized for fixed-size annotations; and
  • differentia width, how many bits should be used per lineage checkpoint to reduce the probability of overestimating relatedness.
Findings are then applied to develop practitioner-oriented guidelines to effectively employ hereditary stratigraphy methodology. Principal takeaway results are,
@@
    tilted retention outperforms steady retention, except in scenarios with very high phylogenetic richness (e.g., drift conditions);
@@
Owing to its inspiration from inference-based phylogenetics work in biology, hereditary stratigraphy is designed to operate in an entirely decentralized manner that is by nature efficient to scale and robust to disruptions or data loss. It is therefore promising to see that reconstruction accuracy of hereditary stratigraphy is also generally robust to scale-up. On the other hand, we found inner node loss — a precision measure — to be sensitive to increases in the number of taxa sampled for reconstruction. This issue, with bit-width differentiae, arises due to increased probability for exactly identical annotations through differentia collision, resulting in clumping of tip nodes into polytomies. It is possible that this problem would be abated in systems with nonsynchronous generations, where tips would be spread apart by generational depth. That said, we did find inner node loss to be largely robust to scale-up of the actual population size of a simulation, with the number of taxa sampled for reconstruction held constant.
@@
Considering a broader perspective on future work, development of hereditary stratigraphy comprises only one aspect of a broader agenda in scaling up digital evolution work. Among other avenues, research will need to explore simulation synchronization schemes (fujimoto1990parallel), best-effort computing approaches (moreno2022best; ackley2020best), emerging hardware architectures (moreno2024trackable; chan2018lenia; heinemann2008artificial), and scalable assays for evolutionary innovation, ecological dynamics, and various forms of complexity (bedau1998classification; dolson2019modes; moreno2024methods; moreno2024case; moreno2024ecology).
@@ Acknowledgement
This research was supported by Michigan State University through the computational resources provided by the Institute for Cyber-Enabled Research and is based upon work supported by the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a Schmidt Futures program. This research was supported in part by funding from the NSF (DEB 1813069). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@
Figure 9: Comparison of reconstruction quality from column- and surface-based annotations. Color coding reflects non-parametric comparison between quality measure values, with red indicating superior surface performance and blue indicating superior column performance. Left column shows tilted retention policies, and right column shows steady retention policies. In cell annotations, +’s indicatve small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 10: Comparison of reconstruction qualities under different differentia retention strategies. Steady retention experiments used column-based implementation; tilted and hybrid retention experiments used surface-based implementation. For heatmap charts, B’s indicate significantly best and w’s significantly worst. Heatmap coloring is nonparametric mean rank among the three algorithms, with blue best and red worst.
@@
Figure 11: Comparison of reconstruction quality from bit- and bite-width differentiae. Color coding reflects non-parametric comparison between quality measure values, with red indicating superior byte-width differentia performance and blue indicating superior bit-width differentia performance. In cell annotations, +’s indicatve small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 12: Comparison of reconstruction quality between small and large downsample and population sizes under steady retention policy. First column considers sample size in isolation, second column considers scaling population size in isolation, and third column considers scaling population and sample size together. Color coding reflects non-parametric comparison between quality measure values, with red indicating degraded reconstruction quality at larger scale and blue indicating improved reconstruction quality at larger scale. Bigger downsample size is 8,000 taxa and smaller downsample size is 500 taxa. Bigger population size is 65,536 and smaller population size is 4,096. Experiments used steady retention policy with column-based implementation. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 13: Comparison of reconstruction quality between small and large downsample and population sizes under hybrid retention policy. First column considers sample size in isolation, second column considers scaling population size in isolation, and third column considers scaling population and sample size together. Color coding reflects non-parametric comparison between quality measure values, with red indicating degraded reconstruction quality at larger scale and blue indicating improved reconstruction quality at larger scale. Bigger downsample size is 8,000 taxa and smaller downsample size is 500 taxa. Bigger population size is 65,536 and smaller population size is 4,096. Experiments used hybrid retention policy with surface-based implementation. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 14: Comparison of reconstruction quality between small and large downsample and population sizes under tilted retention policy. First column considers sample size in isolation, second column considers scaling population size in isolation, and third column considers scaling population and sample size together. Color coding reflects non-parametric comparison between quality measure values, with red indicating degraded reconstruction quality at larger scale and blue indicating improved reconstruction quality at larger scale. Bigger downsample size is 8,000 taxa and smaller downsample size is 500 taxa. Bigger population size is 65,536 and smaller population size is 4,096. Experiments used tilted retention policy with surface-based implementation. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.

### v2

@@ A Practitioner’s Guide to Hereditary Stratigraphy
In simulations employing decentralized parallel and distributed computation, a reconstruction-based approach such as hereditary stratigraphy is more likely to be appropriate. Compared to perfect tracking in this setting, hereditary stratigraphy provides simpler implementation, lower runtime communication overhead, and greater robustness to data loss. Note that if annotation size is not a concern or a low generation count is anticipated, essentially perfect-quality reconstruction can be achieved with hereditary stratigraphy methods. In this case, one could simply use a retention policy that discards no differentiae and a sufficient differentia width to effectively guarantee collisions will not occur (e.g., 32 or 64 bits). However, in most cases, a compromise between phylogeny approximation and per-genome annotation size will be necessary. We cover this in the next section.
@@
In most scenarios, a constant-size annotation will give better runtime performance and ensure fuller use of available memory resources. However, if you need hard guarantees on absolute or recency-proportional inference quality, an annotation size that scales $\mathcal{O}(n)$ or $\mathcal{O}(log(n))$, respectively, with generational depth will be necessary. The hstrat Python package provides fixed_resolution_algo and recency_proportional_resolution_algo policies for such cases (moreno2022hstrat).
@@
Annotation size must compromise between memory-use and communication-bandwidth overhead and quality of reconstructed phylogenies. In cases where annotation size is not a limiting factor, using 256-bit annotation buffers with single-bit differentiae will discern phylogenetic events with about 13% recency-relative precision (tilted) or 1% depth-relative precision (steady) through 1 billion generations. For full-byte differentiae, discussed below, an annotation size on the order of kilobits would be very robust. Where memory use is a limiting factor, however, a 64-bit annotation buffer with single-bit differentiae can give good results.
In picking annotation size, some consideration should be given to the phylogenetic scale of experimental use cases. Triplet distance (accuracy) appears to be largely stable under surveyed increases in population size and number of taxa sampled for reconstruction However, inner node loss can increase when increasing reconstruction sample size. Where this is a concern, annotation size may need to be increased commensurate to intended phylogeny tip count.
Current implementations of surface algorithms are limited to buffer sizes that are even powers of two (i.e., 32, 64, 128, etc.). Where finer gradations are desired, one possibility would be to consider intermediate differentia sizes. For instance, storing 3-bit differentiae over 32 surface sites would occupy 96 bits. Alternatively, alternating depositions across a collection of surfaces might be considered (e.g., a 32-bit tilted surface and a 16-bit steady surface). Note, though, that the latter option would require implementation customizations akin to those used to create the hybrid surface retention policy (moreno2024hsurf). Column algorithms are more flexible in buffer sizing, but in the case of tilted retention, they make less full use of available buffer space.
@@
Hybrid retention policy is a good choice where mild phylodiversity-enhancing factors (ecology, spatial structure, relaxed selection pressure) are expected, or expectations are unclear. Across use cases, hybrid retention often achieves reconstruction quality competitive with the better of steady and tilted policies. Even where it is outperformed by steady or tilted retention, hybrid retention avoids catastrophic failure modes characteristic, in particular, of steady retention. On platforms where steady policy implementation (necessary as a component of hybrid policy) is not readily available, tilted policy should typically suffice. If you do not expect strong phylodiversity-enhancing factors (i.e., no ecology, no spatial structure, high selection pressure), tilted policy can reliably outperform hybrid policy. In rare cases where phylodiversity-enhancing factors are very strong (e.g., pure drift conditions), a steady policy may be appropriate.
@@
Suggested default choice: use bit-size differentiae.
Bit-size differentiae maximize the fraction of correct reconstruction outcomes, but can also introduce incorrect reconstruction outcomes. Byte-size differentiae have very low incorrect reconstruction outcome rates and can reconstruct true polytomies, but have a larger incidence of unresolved reconstruction outcomes (artifactual polytomies). This trade-off results in bit-size differentiae giving more informative reconstructions, as measured by strict triplet distance error. If you need very strong guarantees against incorrect reconstruction outcomes, an even larger differentia size (32 or even 64 bits) may be appropriate.
@@
Another factor that might influence this decision is the software platform targeted. Where packages providing hereditary stratigraphy are not available, surface algorithms are easier to implement owing to only needing to implement one update operation: site selection on the fixed-size surface buffer.
@@
In scenarios where explicitly differentiating between founding clades is paramount, consider adding a systematically assigned founder ID or randomly generated fixed tag. This tag would then be used as a first pass to divvy up end-state annotations into independent sets before feeding them into separate reconstruction processes. When using single-bit differentiae, if several founding lineages persist, spurious collisions can make completely independent clades falsely appear to share some common history.
@@
It is possible to “annotate” differentiae with information about genetic or phenotypic traits associated with corresponding ancestors. Conveniently, because every internal node in a hereditary stratigraphy reconstruction corresponds to a retained differentia, this approach ensures all internal nodes in a reconstructed phylogeny can be associated with a trait value. However, a similar result can be had by saving out sample “fossil” specimens (with corresponding trait information) over the course of a simulation. Because, like extant taxa, these fossils have hereditary stratigraphy annotations, they can readily be incorporated into phylogenetic reconstruction and provide information on ancestral trait states. This approach avoids bloating annotation sizes with dozens of trait values on each genome.
@@
Software implementing column-based approaches is available for Python (moreno2022hstrat) and surface-based approaches are available for Python (moreno2024hsurf) and Zig/Cerebras Software Language (moreno2024wse). Based on community feedback, C/C++ and Rust are priority targets for surface-based implementation ports. However, porting core surface algorithms to another software language should be doable with moderate effort. Surface algorithms are succinct (fewer than 30 lines of code) and accompanying unit tests are provided, making translation to new programming languages straightforward (moreno2024algorithms). We found LLM assistance to be highly effective in translating the surface algorithms from Python to Zig. That said, we encourage interested researchers to reach out if hereditary stratigraphy implementation is not available in their chosen programming language — we would be happy to collaborate in providing implementations in additional languages where there is a use case.
@@
For surface-based annotations, two data components are required: a fixed-size differentia buffer and a generation counter. The differentia store can be implemented as an array of integer data types (e.g., uint8), but for bit differentiae you will likely want to use a raw memory array or an abstraction around it if available (e.g., std::bitset). Differentia store memory should be randomized upon initialization.
@@
When annotations are inherited, they should be updated to add a new differentia and increment the annotation-associated generation counter. Usually, this should be accomplished by adding logic into an existing create offspring routine. For genotype-level tracking (as opposed to individual-level tracking), annotation updates could instead be conducted as a part of an “apply mutation” routine.
@@
For most applications, annotation updates (i.e., appending a new differentia) should correspond directly to generations elapsed. However, in cases where fine-resolution visibility below a certain level is not useful, annotations may be updated less frequently. For tilted policy, in particular, this approach can help preserve ancient differentia coverage for longer.
In circumstances where time-indexed resolution or ultrametricity (tips sharing equal branch lengths from root) is necessary, annotations can be updated on the basis of logical simulation time or even real-time rather than generations elapsed. In this case, annotation updates would be applied, as necessary, on parents’ annotations to bring them up-to-date with the current simulation time whenever a reproduction event occurs. Such an approach may require multiple back-to-back updates if several simulation time clock cycles have elapsed. An alternate approach to achieve a time-calibrated tree would be to attach simulation timestamps to differentiae or to record time-stamped “fossil” taxa, as discussed above.
@@
A typical approach will sample end-state extant agents as taxa for phylogenetic reconstruction. Note that it is not necessary to exhaustively collect the entire end-state population; for many analysis use cases, a representative sample of population members will suffice. Such an approach, however, will not provide information on extinct lineages. Because hereditary stratigraphy reconstruction can stitch together taxa from across widely varying time points, specimens can be sampled during an evolutionary run and be integrated directly into phylogenetic reconstruction with end-state extant genomes. Such intermediate taxa from earlier time points can serve a role akin to “fossils” in natural history. For more advanced analyses, it may be desirable to associate sampled specimens with genetic and/or phenotypic trait data (dolson2019modes; khabbazian2016fast).
@@
To export annotation data for analysis, you will need to save it to file. Workflows that save annotations separately from other genome components using a conventional plain text format (e.g., JSON, CSV) are typically most convenient. The hstrat library provides a number of plug-and-play utilities for serialization of HereditaryStratigraphicColumn objects. Functionality to provide direct support for serializing surface data is a priority item on the hstrat road map. In the meantime, this must be done manually. For each annotation, you will need to store the generation counter and the differentia data. Hex strings provide one potentially convenient approach to encode differentia data. When loading annotation data, you will need to know the retention policy and differentia width used — so make sure these are recorded.
@@ Conclusion
In this work, we have applied empirical annotate-and-reconstruct experiments to benchmark inference quality of hereditary stratigraphy approaches across use cases varying in phylogenetic structure, scale, and allocated annotation space. In these experiments, we consider,
  • differentia retention: whether annotation space should be allocated for finer resolution in discerning recent phylogenetic events,
  • annotation implementation: comparing existing column-based approaches to newer surface-based approaches optimized for fixed-size annotations, and
  • differentia width: how many bits should be used per lineage checkpoint to reduce the probability of spuriously overestimating relatedness.
Findings are then applied to develop practitioner-oriented guidelines to effectively employ hereditary stratigraphy methodology. Principal results are,
@@
    tilted retention produces better reconstruction quality than steady retention, except in scenarios with very high phylogenetic richness (e.g., drift conditions);
@@
Owing to its inspiration from inference-based phylogenetics work in biology, hereditary stratigraphy is designed to operate in an entirely decentralized manner that is, by nature, efficient to scale and robust to disruptions or data loss. It is therefore promising to see that reconstruction accuracy of hereditary stratigraphy is also generally robust to scale-up. On the other hand, we found inner node loss — a precision measure — to be sensitive to increases in the number of taxa sampled for reconstruction. This issue, with bit-width differentiae, arises due to increased probability for exactly identical annotations through differentia collision, resulting in clumping of tip nodes into polytomies. This problem may be abated in systems with nonsynchronous generations, where tips are spread apart by generational depth. That said, we did find inner node loss to be largely robust to scale-up of the actual population size of a simulation, with the number of taxa sampled for reconstruction held constant.
@@
Considering a broader perspective on future work, development of hereditary stratigraphy comprises only one aspect of a broader agenda in scaling up digital evolution experiments. Among other avenues, research will need to explore simulation synchronization schemes (fujimoto1990parallel), best-effort computing approaches (moreno2022best; ackley2020best), emerging hardware architectures (moreno2024trackable; chan2018lenia; heinemann2008artificial), and scalable assays for evolutionary innovation, ecological dynamics, and various forms of complexity (bedau1998classification; dolson2019modes; moreno2024methods; moreno2024case; moreno2024ecology).
@@ Acknowledgement
This research was supported by Michigan State University through the computational resources provided by the Institute for Cyber-Enabled Research and is based upon work supported by the Eric and Wendy Schmidt AI in Science Postdoctoral Fellowship, a Schmidt Sciences program. This research was supported in part by funding from the NSF (DEB 1813069). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@
Figure 9: Comparison of reconstruction quality from column- and surface-based annotations. Color coding reflects non-parametric comparison between quality measure values, with red indicating superior surface performance and blue indicating superior column performance. Left column shows tilted retention policies, and right column shows steady retention policies. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 10: Comparison of reconstruction quality across differentia retention strategies. Steady retention experiments used column-based implementation; tilted and hybrid retention experiments used surface-based implementation. For heatmap charts, B’s indicate significantly best and w’s significantly worst. Heatmap coloring shows nonparametric mean rank among the three algorithms, with blue best and red worst.
@@
Figure 11: Comparison of reconstruction quality from bit- and byte-width differentiae. Color coding reflects non-parametric comparison between quality measure values, with red indicating superior byte-width differentia performance and blue indicating superior bit-width differentia performance. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 12: Comparison of reconstruction quality between small and large downsample and population sizes under steady retention policy. First column considers sample size in isolation, second column considers scaling population size in isolation, and third column considers scaling population and sample size together. Color coding reflects non-parametric comparison between reconstruction quality measure values, with red indicating degraded reconstruction quality at larger scale and blue indicating improved reconstruction quality at larger scale. Larger downsample size is 8,000 taxa and smaller downsample size is 500 taxa. Larger population size is 65,536 and smaller population size is 4,096. Experiments used steady retention policy with column-based implementation. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 13: Comparison of reconstruction quality between small and large downsample and population sizes under hybrid retention policy. First column considers sample size in isolation, second column considers scaling population size in isolation, and third column considers scaling population and sample size together. Color coding reflects non-parametric comparison between reconstruction quality measure values, with red indicating degraded reconstruction quality at larger scale and blue indicating improved reconstruction quality at larger scale. Larger downsample size is 8,000 taxa and smaller downsample size is 500 taxa. Larger population size is 65,536 and smaller population size is 4,096. Experiments used hybrid retention policy with surface-based implementation. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.
@@
Figure 14: Comparison of reconstruction quality between small and large downsample and population sizes under tilted retention policy. First column considers sample size in isolation, second column considers scaling population size in isolation, and third column considers scaling population and sample size together. Color coding reflects non-parametric comparison between reconstruction quality measure values, with red indicating degraded reconstruction quality at larger scale and blue indicating improved reconstruction quality at larger scale. Larger downsample size is 8,000 taxa and smaller downsample size is 500 taxa. Larger population size is 65,536 and smaller population size is 4,096. Experiments used tilted retention policy with surface-based implementation. In cell annotations, +’s indicate small, medium, and large effect sizes using the Cliff’s delta statistic and *’s indicate statistical significance at $\alpha=0.05$ via Mann-Whitney U test.